OpenStudy (anonymous):

HELP!!!!!!!!!!!! FAST!!!!!!!!!!! For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 113x + 4y = 24

OpenStudy (anonymous):

2x + y = 11 3x + 4y = 24 Solve the first one for y

OpenStudy (anonymous):

2x + y = 11 y = 11 - 2x now plug that into the place of y, and then solve for x. c:

OpenStudy (anonymous):

x-cost of aples y-cost of oranges Two apples + 1 orange cost 11 3 apples + 4 oranges cost 24

OpenStudy (anonymous):

Jus a sec. like this

OpenStudy (anonymous):

2x + y = 11 3x + 4y = 24 take equation 1 and substitute into 2: y = 9 - 2x 3x + 4(9 - 2x) = 26 3x + 36 - 8x = 26 -5x = -10 x = 2 y = 9 - 4 y = 5 (2,5)